Opened 11 years ago

Last modified 7 years ago

#160 assigned flight processing

GB08/08, flight day 134b/2008, wytham woods

Reported by: mark1 Owned by: adbe
Priority: alpha 4 high Milestone: Reprocessing
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by mark1)

Data location: ~arsf/arsf_data/in_progress/2008/flight_data/uk/GB08_08-2008_134b_wytham_woods

Data arrived from ARSF via initial SATA disk transfer on 20th May 2008.

Scientific details:

PI: R Harding

Sensors:

  • ATM
  • Eagle
  • Hawk

Quicklook

TBD.

Change History (13)

comment:1 Changed 11 years ago by mark1

Have removed ATM files 001 - 007 since these relate to project 134a, GB08/18. They can be found in the directory for GB08/18 if required.

comment:2 Changed 11 years ago by mark1

Project code on logsheet is (probably!) wrong. Sticking with GB08/05 since this was project code for previous wytham woods data set. Gary says they will all have same proj code. Also P.I.'s are different on each wytham woods project logsheet...Gary say's send all raw data to Andrew Wilson with applanix.

comment:3 Changed 11 years ago by mark1

Starting to think project code is wrong and should be GB08/08 not GB08/05

comment:4 Changed 11 years ago by mark1

  • Description modified (diff)
  • Summary changed from GB08/05, flight day 134b/2008, wytham woods to GB08/08, flight day 134b/2008, wytham woods

After phone conversation with gary on 2/6/08 have removed all reference to GB08/05 and replaced with GB08/08. PI is R Harding although all data will go to A Wilson.

comment:5 Changed 11 years ago by mggr

Applanix processed and synced to ~arsf directory. Should be ready to go (albeit untested).

comment:6 Changed 11 years ago by mark1

Project has been copied onto hard disk in preparation for delivery.

comment:7 Changed 11 years ago by mark1

Data has been delivered on the 7th July 2008

comment:8 Changed 10 years ago by mggr

  • Resolution set to fixed
  • Status changed from new to closed

Data sent to NEODC for archiving.

comment:9 Changed 7 years ago by benj

  • Resolution fixed deleted
  • Status changed from closed to reopened

Level 1 data doesn't seem to have gotten from AKW to NEODC. Eagle/Hawk need to be reprocessed and sent to NEODC, please notify Charles George (ctg@…) when done.

comment:10 Changed 7 years ago by benj

  • Owner set to adbe
  • Status changed from reopened to assigned

comment:11 Changed 7 years ago by adbe

Reprocessed navigation. The sbet did not cover the first line for eagle and hawk.

comment:12 Changed 7 years ago by adbe

Created Delivery

comment:13 Changed 7 years ago by adbe

Ready to deliver at:

/users/rsg/arsf/workspace/GB08_08-2008_134b_wytham_woods/processing/delivery/GB08_08-134b-hyperspectral-20120201


Note: See TracTickets for help on using tickets.